Circular No. 8160

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

NOVA IN THE LARGE MAGELLANIC CLOUD On June 19, W. Liller (Viña del Mar, Chile) reported his discovery

of a possible nova (mag 11.0) in the LMC from unfiltered Technical Pan film exposed on June 17.974 and 17.983 UT, noting that nothing brighter than mag 14 had been detected on exposures from June 2.970. An unfiltered CCD exposure taken by L. A. G. Monard (Pretoria, S. Africa) on June 20.167 yielded mag 11.4 and precise position α = 5^{h}08^{m}25^{s}.59, δ = −68^{o}26′22′′.3 (equinox 2000.0), and he noted that is close to the location of a UCAC star of mag 13.8 at position end figures 25^{s}.46, 22′′.3, and a USNO-A2.0 star (blue mag 16.9, red mag 14.6) at position end figures 25^{s}.42, 22′′.4. CCD V magnitudes from Liller: June 20.958, 12.2 (through clouds); 21.958, 12.4; 23.959, 12.6. Available visual magnitudes by A. Pearce, Nedlands, W. Australia: June 20.458, 12.3; 20.858, 12.1; 25.455, 13.0. H. E. Bond, Space Telescope Science Institute, reports that S. Gonzalez obtained several blueward spectrograms on June 30 with the Cerro Tololo 1.5-m reflector. Bond notes that the star shows very broad Balmer emission lines, so there is little doubt that it is a classical nova.

η CARINAE

P. A. Whitelock, F. Marang, and L. Crause, South African Astronomical Observatory, report that observations show that the anticipated fading of η Car (cf. IAUC 8154) at infrared wavelengths (JHKL) started between June 19 (L-band) and June 24 (J-band); following Feast et al. (2001, MNRAS 322, 742), infrared minimum can be expected around July 13.

Further to IAUC 8154, M. F. Corcoran writes that an RXTE observation of η Car on June 29 shows that the 2–10-keV x-ray emission has fallen to 8.3 × 10^{−13} erg s^{−1} cm^{−2}; η Car has not been this faint in x-rays since the last x-ray minimum during Dec. 1997–Jan. 1998. Information on known space and ground-based observations is given at http://lheawww.gsfc.

nasa.gov/users/corcoran/eta car/2003.5/.

V4745 SAGITTARII Regarding the increase in brightness reported on IAUC 8155, A. C.

Gilmore reports additional photometry of V4745 Sgr by P. M. Kilmartin at Mt. John Observatory (cf. IAUC 8123): June 25.620 UT, V = 8.96, U − B = −0.98, B − V = +0.12, V − R = +0.27, V − I = +0.38, airmass 1.08 (uncertainties < 0.02 mag in all colors).

2003 June 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green